# Language Fusion for Parameter-Efficient Cross-lingual Transfer

### v1

@@ 1 Introduction
In this work, we introduce Fusion for Language Representations (FLARE), a novel approach that merges latent representations from different languages within lower-dimensional adapter bottlenecks to enable parameter-efficient XLT. By merging representations from high-resource languages like English into target language representations through lightweight fusion functions, such as addition or multiplication, FLARE facilitates effective cross-lingual information transfer with minimal computational overhead. As illustrated in Figure 1, FLARE performs token-wise fusion of source and target language representations within each transformer block, without adding additional parameters to LoRA and maintaining computational efficiency.
Our experiments demonstrate FLARE’s versatility and effectiveness across tasks like natural language inference, sentiment classification, and question answering, using encoder-only, encoder-decoder, and decoder-only multilingual pre-trained language models (mPLMs). It is particularly beneficial for downstream tasks that involve text generation, such as question answering. For instance, FLARE improves the exact match performance for Llama 3.1 and Gemma 2 on the TyDiQA dataset by 4.9% and 2.2%, respectively. Further experiments illustrate that computational efficiency can be further enhanced by using latent translations as source language inputs in FLARE, and demonstrate the flexibility of the method, which is orthogonal to the choice of mPLMs and MT systems.
Contributions. 1) We introduce FLARE, a novel method that fuses language representations within adapter bottlenecks for parameter-efficient cross-lingual transfer. 2) Our approach improves performance across diverse multilingual downstream tasks, particularly benefiting text generation tasks (e.g., applied to question answering). 3) We demonstrate the adaptability of our approach by incorporating machine translation encoder representations directly into the mPLM.
@@ 2 Related Work
Representation fusion has also been applied to integrate information across different modalities, such as vision and language (Fang et al., 2021; Ramnath et al., 2021). For instance, Qu et al. (2025) used feature routing in cross-modal vision-language tasks, guiding language model representations through LoRA bottlenecks using the last hidden state of a vision model. Our work differs in its scope and fusion methodology: FLARE extracts significantly richer representations from the source and target languages by capturing layer-wise representations for each transformer block in the mPLMs. Moreover, by ensuring dimensional alignment, we perform token-wise representation fusion within adapter bottlenecks, thereby transferring finer-grained information across languages.
@@ 3.1 Language Representation Fusion
Our methodology is based on the hypothesis that incorporating English with target language representations enhances cross-lingual knowledge transfer and distills task-relevant information into the target language. We assume (typically MT-created) parallel corpora during task fine-tuning, where are instances in the respective source and target language. Our methodology particularly focuses on employing machine-translated ‘silver’ parallel data, akin to translate-train and translate-test settings, as we believe this approach is the most realistic in practice.
@@
To address these limitations, we propose FLARE, a method for representation-level language fusion within bottleneck adapters, as illustrated in Figure 1. Instead of extending the input, FLARE processes source and target language representations independently and fuses them only within the adapters, thus preserving computational efficiency. Source language representations , extracted from the frozen mPLM without adapters, and target language representation at transformer block are down-projected using and combined with fusion function (see Section 3.2) to create a fused representation , where with sequence length and bottleneck dimensions . We utilize the source representation , which has been processed by the subsequent transformer block, to leverage task-specific information extracted from the source language. Following a standard LoRA procedure, this fused low-rank representation is then up-projected and added to the frozen attention outputs to form the target language output representation of the attention block.
@@ 3.2 Fusion Functions
cross-attention:333Although cross-attention modules add parameters to the adapters, the low bottleneck dimensions , typically smaller than 64, minimize the parameter count in comparison to the model’s internal dimensions . Specifically, we utilize a single cross-attention head to maintain efficiency. , and are the weight matrices of the query, key and value projections in the adapter , respectively, and ′ denotes the matrix transpose. We focus on lightweight linear transformations to maintain parameter and computational efficiency.
@@ 4.1 Underlying Models and Baselines
Fine-Tuning Setup. We follow a modular XLT approach where the mPLM is fine-tuned on English task data and subsequently adapted using task data in the target language (Zhao et al., 2021). For decoder-only models like Llama and Gemma, we use a causal language modeling objective for fine-tuning, and the models generate predictions as text accordingly. We employ the QLoRA fine-tuning approach with 4-bit quantization and insert LoRA adapters in all linear layers for Llama and Gemma models (Dettmers et al., 2023). Importantly, we apply representation fusion in FLARE only in the attention modules, ensuring a consistent experimental setup across different transformer architectures. The LoRA configurations use and , and the hyperparameter configurations for each model are detailed in Table 5 in the appendix.
Baselines. We evaluate FLARE against several baselines, including zero-shot cross-lingual transfer, translate-test, as well as translate-train methods such as regular LoRA fine-tuning, X-Mixup, and input-level fusion. All translate-train models are trained with the same LoRA configurations. Unless otherwise specified, FLARE models are trained using the add+relu fusion function, with a detailed comparison of fusion functions presented in Table 2. Model checkpoints are selected based on validation data that was machine-translated from English to the respective target languages.
@@ 4.3 Machine Translations
Model | XNLI | TyDiQA | NusaX | Avg. |
Zero-Shot Cross-Lingual Transfer (models are trained on English data) | ||||
XLM-R Large | ||||
mT5-XL | ||||
Llama 3.1 8B | ||||
Gemma 2 9B | ||||
Translate-Test (test data is translated to English) | ||||
XLM-R Large | ||||
mT5-XL | ||||
Llama 3.1 8B | ||||
Gemma 2 9B | ||||
XLM-R Large w/ LoRA | ||||
w/ X-Mixup | ||||
w/ input-level fusion | ||||
w/ FLARE MT | ||||
w/ FLARE | ||||
mT5-XL w/ LoRA | ||||
w/ X-Mixup | ||||
w/ input-level fusion | ||||
w/ FLARE MT | ||||
w/ FLARE | ||||
Llama 3.1 8B w/ LoRA | ||||
w/ X-Mixup | ||||
w/ input-level fusion | ||||
w/ FLARE MT | ||||
w/ FLARE | ||||
Gemma 2 9B w/ LoRA | ||||
w/ X-Mixup | ||||
w/ input-level fusion | ||||
w/ FLARE MT | ||||
@@ 5 Results and Discussion
Main Results. The results displayed in Table 1 confirm our hypothesis that task-specific knowledge can be efficiently transferred from English to other languages within adapter bottlenecks. Our proposed approach, FLARE, consistently surpasses all baselines across various tasks, demonstrating robust performance and validating the effectiveness of our method. It improves the average performance, averaged across metrics for all tasks, by 2.14% and 1.27% for Llama and Gemma, respectively, when compared to standard LoRA fine-tuning. The most substantial performance gains are observed on the TyDiQA dataset, particularly for text generation tasks with decoder-only models. FLARE significantly improves performance on this dataset, with the largest gains achieved on Indonesian, Russian, and Swahili. This suggests that latent representation fusion with FLARE works best for text generation when the target languages have a similar word order to the source language, in this case, subject-verb-object. However, we also observe substantial performance gains for the Llama model on Telugu, which has a different word order than English, indicating that FLARE can still achieve considerable improvements even when the word order differs.
@@
For encoder-only models like XLM-R Large and encoder-decoder models like mT5, latent MT fusion provides notable performance benefits compared to standard LoRA fine-tuning, X-Mixup, and input-level fusion, as shown in Table 2. However, for decoder-only models like Llama and Gemma, we do not observe significant performance benefits from latent MT fusion. This suggests that reducing the computational resources for processing the source language representations in regular FLARE can negatively impact cross-lingual transfer performance, particularly for larger models. Nonetheless, it provides a resource-efficient alternative to regular FLARE for smaller mPLMs by avoiding the need for decoding in the MT and eliminating the forward pass for the source language representations.
@@
Impact of Adapter Capacity. Our ablation study, presented in Table 3, investigates the impact of adapter capacity on FLARE’s performance. The results reveal that small bottleneck sizes () yield optimal performance for XLM-R Large on the TyDiQA and NusaX datasets. This finding is consistent with the observations in the original LoRA paper (Hu et al., 2022), indicating that the introduction of our fusion adapter does not affect the intrinsic rank of the tasks.
@@ 6 Conclusion
In this paper, we introduced Fusion for Language Representations (FLARE), a parameter-efficient method for cross-lingual transfer (XLT) that enhances representation quality and downstream performance for languages other than English. Our experimental results demonstrate that FLARE consistently outperforms strong XLT baselines, including target language fine-tuning with LoRA adapters, X-Mixup, and input-level fusion, on various natural language understanding tasks. FLARE demonstrates robust performance, even for lower-quality machine translations. A key takeaway is that FLARE remains more parameter-efficient compared to benchmarked baseline approaches, while yielding superior performance. Furthermore, FLARE provides most substantial performance benefits for multilingual questions answering with decoder-only language models such as those from Llama and Gemma model families.
@@ Appendix A Detailed Evaluation Results
Table 4 presents the performance of FLARE and input-level fusion when using gold translations for fusion, as opposed to machine translations generated by NLLB. The results demonstrate that input-level fusion performance is sensitive to the quality of English input provided. Notably, when gold translations are available, input-level fusion replicates English performance, indicating that it heavily relies on the quality of English inputs. In contrast, FLARE balances the fusion of source and target language information, as evident from the findings in Figure 5. While input-level fusion outperforms FLARE when gold translations are available, FLARE achieves significantly higher performance in the more realistic setting using machine-translated data.
@@ Appendix E Another Ablation: Representation Fusion during Training Only
To investigate the importance of utilizing source language representations during inference, we modified FLARE to restrict representation fusion to the training phase only. Specifically, we limited the fusion with source language representations to 50% of the training instances and excluded source language data during inference. This evaluates cross-lingual transfer capabilities based on instance-independent patterns learned from source language representations during training. Our findings reveal that fusion adapters struggle to learn patterns that are independent of specific instances from source language representations during training. As a result, when implemented in the XLM-R Large model on the NusaX test set, the performance of the train-only FLARE variant decreased by 30%. Crucially, this significant drop underscores the importance of incorporating source language representations during inference to achieve effective cross-lingual adaptation.

### v2

@@ 1 Introduction
In this work, we introduce Fusion for Language Representations (FLARE), a novel approach that merges latent representations from different languages within lower-dimensional adapter bottlenecks to enable parameter-efficient XLT. By merging representations from high-resource languages like English into target language representations through lightweight fusion functions, such as addition or multiplication, FLARE facilitates effective cross-lingual information transfer with minimal computational overhead. As illustrated in the schematic overview in Figure 1, FLARE performs token-wise fusion of source and target language representations within each transformer block, without adding additional parameters to LoRA and maintaining computational efficiency. FLARE is applied during task-specific fine-tuning, enabling models to extract knowledge from both the source and target languages and apply it effectively in the target language.
Our experiments demonstrate FLARE’s versatility and effectiveness across tasks like natural language inference, sentiment classification, and question answering, using encoder-only, encoder-decoder, and decoder-only multilingual pre-trained language models (mPLMs). It is particularly beneficial for downstream tasks that involve text generation, such as question answering. For instance, FLARE improves the absolute exact match performance for Llama 3.1 and Gemma 2 on the TyDiQA dataset by 4.9% and 2.2%, respectively. Further experiments illustrate that computational efficiency can be further enhanced by using latent translations as source language inputs in FLARE, and demonstrate the flexibility of the method, which is orthogonal to the choice of mPLMs and MT systems.
Contributions.
1) We introduce FLARE, a novel method that fuses language representations within adapter bottlenecks for parameter-efficient cross-lingual transfer. 2) Our approach improves performance across diverse multilingual downstream tasks, particularly benefiting text generation tasks (e.g., applied to question answering). 3) We demonstrate the adaptability of our approach by incorporating machine translation encoder representations directly into the mPLM.
@@ 2 Related Work
Representation fusion has also been applied to integrate information across different modalities, such as vision and language (Fang et al., 2021; Ramnath et al., 2021). For instance, Qu et al. (2025) used feature routing in cross-modal vision-language tasks, guiding language model representations through LoRA bottlenecks using the last hidden state of a vision model. Our work differs in its scope and fusion methodology: FLARE extracts richer representations from the source and target languages by capturing layer-wise representations for each transformer block in the mPLMs. Moreover, by ensuring dimensional alignment, we perform token-wise representation fusion within adapter bottlenecks, thereby transferring finer-grained information across languages.
@@ 3.1 Language Representation Fusion
Our methodology is based on the hypothesis that incorporating English with target language representations enhances cross-lingual knowledge transfer and distills task-relevant information into the target language. We assume (typically MT-created) parallel corpora during task fine-tuning, where are instances in the respective source and target language. Our methodology focuses on using machine-translated parallel data, where widely available English data is translated into the target language for training, and target language data is translated into English during inference. This approach reflects practical real-world scenarios, as human-annotated data is often scarce and costly.
@@
To address these limitations, we propose FLARE, a method for representation-level language fusion within bottleneck adapters, as illustrated in Figure 2. Instead of extending the input, FLARE processes source and target language representations independently and fuses them only within the adapters, thus preserving computational efficiency. Source language representations , extracted from the frozen mPLM without adapters, and target language representation at transformer block are down-projected using and combined with fusion function (see Section 3.2) to create a fused representation , where with sequence length and bottleneck dimensions (Qu et al., 2025). We utilize the source representation , which has been processed by the subsequent transformer block, to leverage task-specific information extracted from the source language. Following a standard LoRA procedure, this fused low-rank representation is then up-projected and added to the frozen attention outputs to form the target language output representation of the attention block.
@@ 3.2 Fusion Functions
cross-attention:333Although cross-attention modules add parameters to the adapters, the low bottleneck dimensions , typically smaller than 64, minimize the parameter count in comparison to the model’s internal dimensions . Specifically, we utilize a single cross-attention head to maintain efficiency.
, and are the weight matrices of the query, key and value projections in the adapter , respectively, and ′ denotes the matrix transpose. We focus on lightweight linear transformations to maintain parameter and computational efficiency.
@@ 4.1 Underlying Models and Baselines
Fine-Tuning Setup. We follow a modular XLT approach where the mPLM is fine-tuned on English task data and subsequently adapted using task data in the target language (Zhao et al., 2021). For decoder-only models like Llama and Gemma, we use a causal language modeling objective for fine-tuning, and the models generate predictions as text accordingly. We employ the QLoRA fine-tuning approach with 4-bit quantization and insert LoRA adapters in all linear layers for Llama and Gemma models (Dettmers et al., 2023). Importantly, we apply representation fusion in FLARE only in the attention modules, ensuring a consistent experimental setup across different transformer architectures. The LoRA configurations use and , where corresponds to the bottleneck dimensions and is a scaling factor applied to the low-rank updates before merging them with the frozen attention outputs. Detailed hyperparameter configurations for each model can be found in Table 5 in the appendix.
Baselines. We evaluate FLARE against several baselines, including zero-shot cross-lingual transfer, translate-test, as well as translate-train methods such as regular LoRA fine-tuning, X-Mixup, and input-level fusion. We focus on baseline methods that fine-tune on task-specific data in the target language, as these provide the most competitive performance. All translate-train models are trained with the same LoRA configurations. Unless otherwise specified, FLARE models are trained using the add+relu fusion function, with a detailed comparison of fusion functions presented in Table 2. Model checkpoints are selected based on validation data that was machine-translated from English to the respective target languages.
@@ 4.3 Machine Translations
Model | XNLI | TyDiQA | NusaX | Avg. | Avg. rank |
Zero-Shot Cross-Lingual Transfer (models are trained on English data) | |||||
XLM-R Large | |||||
mT5-XL | |||||
Llama 3.1 8B | |||||
Gemma 2 9B | |||||
Translate-Test (test data is translated to English) | |||||
XLM-R Large | |||||
mT5-XL | |||||
Llama 3.1 8B | |||||
Gemma 2 9B | |||||
Translate-Train (models are trained on training data translated to the target language) | |||||
XLM-R Large w/ LoRA | |||||
w/ X-Mixup (Yang et al., 2022) | |||||
w/ input-level fusion | |||||
w/ FLARE MT | |||||
w/ FLARE | |||||
mT5-XL w/ LoRA | |||||
w/ X-Mixup (Yang et al., 2022) | |||||
w/ input-level fusion | |||||
w/ FLARE MT | |||||
w/ FLARE | |||||
Llama 3.1 8B w/ LoRA | |||||
w/ X-Mixup (Yang et al., 2022) | |||||
w/ input-level fusion | |||||
w/ FLARE MT | |||||
w/ FLARE | |||||
Gemma 2 9B w/ LoRA | |||||
w/ X-Mixup (Yang et al., 2022) | |||||
w/ input-level fusion | |||||
w/ FLARE MT | |||||
@@ 5 Results and Discussion
Main Results. The results displayed in Table 1 confirm our hypothesis that task-specific knowledge can be efficiently transferred from English to other languages within adapter bottlenecks. Our proposed approach, FLARE, consistently surpasses all baselines across various tasks, demonstrating robust performance and validating the effectiveness of our method. It improves the average performance, averaged across metrics for all tasks, by 2.14% and 1.27% for Llama and Gemma, respectively, when compared to standard LoRA fine-tuning. The most substantial performance gains are observed on the TyDiQA dataset, particularly for text generation tasks with decoder-only models. FLARE substantially improves performance on this dataset, with the largest gains achieved on Indonesian, Russian, and Swahili. This suggests that latent representation fusion with FLARE works best for text generation when the target languages have a similar word order to the source language, in this case, subject-verb-object. However, we also observe substantial performance gains for the Llama model on Telugu, which has a different word order than English, indicating that FLARE can still achieve considerable improvements even when the word order differs.
@@
For encoder-only models like XLM-R Large and encoder-decoder models like mT5, latent MT fusion provides notable performance benefits compared to standard LoRA fine-tuning, X-Mixup, and input-level fusion, as shown in Table 2. However, for decoder-only models like Llama and Gemma, we do not observe substantial performance benefits from latent MT fusion. This suggests that reducing the computational resources for processing the source language representations in regular FLARE can negatively impact cross-lingual transfer performance, particularly for larger models. Nonetheless, it provides a resource-efficient alternative to regular FLARE for smaller mPLMs by avoiding the need for decoding in the MT and eliminating the forward pass for the source language representations.
@@
Impact of Adapter Capacity. We investigate the impact of the adapter capacity on FLARE’s performance, with results presented in Table 3. The findings indicate that small bottleneck sizes () yield optimal performance for XLM-R Large on the TyDiQA and NusaX datasets. This observation is consistent with the findings in the original LoRA paper (Hu et al., 2022), indicating that the introduction of our fusion adapter does not affect the intrinsic rank of the tasks.
@@ 6 Conclusion
In this paper, we introduced Fusion for Language Representations (FLARE), a parameter-efficient method for cross-lingual transfer (XLT) that enhances representation quality and downstream performance for languages other than English. Our experimental results demonstrate that FLARE consistently outperforms strong XLT baselines, including target language fine-tuning with LoRA adapters, X-Mixup, and input-level fusion, on various natural language understanding tasks. FLARE demonstrates robust performance, even for lower-quality machine translations. A key takeaway is that FLARE remains more parameter-efficient compared to benchmarked baseline approaches, while yielding superior performance. Furthermore, FLARE provides most substantial performance benefits for multilingual question answering with decoder-only language models such as those from Llama and Gemma model families.
@@ 8 Limitations
The reliance on machine translation also pose challenges for extremely low-resource languages, particularly where translation quality suffers or where cultural nuances are difficult to preserve (Alemayehu et al., 2024; Conia et al., 2024). To examine the impact of translation quality on FLARE, we conducted a manual error analysis on 200 randomly selected TyDiQA examples using Llama 3.1 and Gemma 2. We found no cases where machine translation quality was responsible for model errors. Most errors were due to answers unrelated to the question, while a smaller subset consisted of correct answers that did not match the dataset’s ground truth. Additionally, the flexibility of FLARE’s fusion functions provides opportunities to incorporate cultural context more explicitly in future work.
@@ Appendix A Detailed Evaluation Results
Table 4 presents the performance of FLARE and input-level fusion when using gold translations for fusion, as opposed to machine translations generated by NLLB. The results demonstrate that input-level fusion performance is sensitive to the quality of English input provided. Notably, when gold translations are available, input-level fusion replicates English performance, indicating that it heavily relies on the quality of English inputs. In contrast, FLARE balances the fusion of source and target language information, as evident from the findings in Figure 6. While input-level fusion outperforms FLARE when gold translations are available, FLARE achieves substantially higher performance in the more realistic setting using machine-translated data.
@@
Table 9 reports p-values from the Pitman permutation tests comparing FLARE and FLARE MT to baseline methods, based on average per-language performance from Tables 6, 7, and 8. The results indicate that FLARE and FLARE MT yield consistent and often statistically significant improvements over regular LoRA, X-Mixup, and input-level fusion, with particularly strong gains for the decoder-only Llama and Gemma models.
@@ Appendix E Another Ablation: Representation Fusion during Training Only
To investigate the importance of utilizing source language representations during inference, we modified FLARE to restrict representation fusion to the training phase only. Specifically, we limited the fusion with source language representations to 50% of the training instances and excluded source language data during inference. This evaluates cross-lingual transfer capabilities based on instance-independent patterns learned from source language representations during training. Our findings reveal that fusion adapters struggle to learn patterns that are independent of specific instances from source language representations during training. As a result, when implemented in the XLM-R Large model on the NusaX test set, the performance of the train-only FLARE variant decreased by 30%. Crucially, this drop in performance underscores the importance of incorporating source language representations during inference to achieve effective cross-lingual adaptation.
@@
Dataset | Model | LoRA | X-Mixup | Input-level fusion |
XNLI | XLM-R w/ FLARE MT | *** | *** | *** |
w/ FLARE | *** | *** | ||
mT5-XL w/ FLARE MT | *** | *** | *** | |
w/ FLARE | *** | *** | *** | |
Llama 3.1 8B w/ FLARE MT | *** | |||
w/ FLARE | *** | |||
Gemma 2 9B w/ FLARE MT | *** | *** | ||
w/ FLARE | *** | *** | ||
TyDiQA | XLM-R w/ FLARE MT | |||
w/ FLARE | ** | |||
mT5-XL w/ FLARE MT | ||||
w/ FLARE | ||||
Llama 3.1 8B w/ FLARE MT | ** | *** | *** | |
w/ FLARE | *** | *** | ||
Gemma 2 9B w/ FLARE MT | ** | |||
w/ FLARE | ** | *** | ||
NusaX | XLM-R w/ FLARE MT | |||
w/ FLARE | ** | *** | ||
mT5-XL w/ FLARE MT | ** | ** | ||
w/ FLARE | ** | |||
Llama 3.1 8B w/ FLARE MT | * | |||
w/ FLARE | ** | |||
Gemma 2 9B w/ FLARE MT | * | ** | ||
w/ FLARE | ** | * |